The accompanying data on x = head circumference z score (a comparison score with peers of the same age - a positive score suggests a larger size than for peers) at age 6 to 14 months and y = volume of cerebral grey matter (in ml) at age 2 to 5 years were read from a graph in an article.

Cerebral Grey
Matter (ml) 2-5 yr
Head Circumfer-
ence z Scores at
6-14 Months
680 -0.72
690 1.6
700 -0.27
720 0.28
740 0.33
740 1.9
750 1.5
750 2.4
760 1.5
780 1.5
790 2.4
810 2.5
815 3.2
820 2.6
825 0.93
835 2.75
840 2.7
845 2.6

(a) What is the value of the correlation coefficient? (Give the answer to two decimal places.)
r =


(b) Find the equation of the least-squares line. (Give the answer to two decimal places.)
=




(c) Predict the volume of cerebral grey matter for a child whose head circumference z score at age 12 months was 1.8. (Give the answer to two decimal places.)
Answer #1

EXCEL (Data Analysis) Procedure: Step 1: In EXCEL sheet, enter Head Circumfer-z Scores at 6-14 Months in one column and enterSUMMARY OUTPUT Regression Statistics Multiple R R Square Adjusted R Square Standard Error Observations 0.755733 0.571132 0.54(c) û = 712.77 +35.70 x =712.77 +35.70(1.8) = 712.77 +64.26 = 777.03Thus the best predicted volume of cerebral grey matter for a child whose head circumference z score at age 12 months was 1.8